OpenStudy (anthony0223990):

An artist wants to make alabaster pyramids using a block of alabaster with a volume of 576 cubic inches. She plans to make each pyramid with a square base area of 3 square inches and a height of 4 inches. At most, how many pyramids can the artist make from the block of alabaster?

OpenStudy (anthony0223990):

@DrummerGirl3

OpenStudy (drummergirl3):

V = 1/3 * area of base * height V = 1/3 * 3 * 4 V = 4 in^3 576/4 = 144 => max number of pyramids made out of the block of alabaster. b) => answer mi amigo